A livestock company reports that the mean weight of a group of young steers is 1110 pounds with a standard deviation of 61 popls answer question a,b,c
Answer #1


W= Jllo pounds 6= 61 pounds (a) for P(x) 1150) PC xJITS0) = |- P(x5 1956) =1-P( 23. 1150- 1110 19 - 1-P (254) = 1-P(Z = 0.655